4xyz=pq+2px^2y+2qxy^2


1
Expert's answer
2021-03-12T06:02:09-0500

"4xyz = pq + 2px^2y + 2qxy^2\\
f(x,y,z,p,q) = pq + 2px^2y + 2qxy^2 - 4xyz = 0 ... (*)\\ f_x = 4pxy+2qy^2 - 4yz \\ f_y = 2px^2 + 4 qxy - 4xz\\ f_z = -4xy \\ f_p = q+2x^2y\\ f_q =p+2xy^2"


By charpit method,

"\dfrac{dx}{-q-2x^2y} = \dfrac{dy}{-p-2xy^2} = \dfrac{dz}{-2pq-2px^2y - 2qxy^2} = \dfrac{dp}{2qy^2-4yz} = \dfrac{dq}{2px^2-4xz}"


"\dfrac{xdy-ydx}{qy-px} = \dfrac{xdp-ydq}{2xy(qy-px)} \\ xdp -ydq = 0\\ xp = yq \\ p = \frac{y}{x}q"

Substituting into (*), we have

"yq^2 +4x^2y^2q-4x^2yz = 0 \\
q = -2x^2y \pm 2x \sqrt{x^2y^2+z} \\"

Taking positive sign only

"q = -2x^2y + 2x \sqrt{x^2y^2 +z} \\ \therefore p = -2xy^2 + 2y \sqrt{x^2y^2+z}"

Substituting the value of p and q into pdx+qdy = dz and also integrating gives

"(-2xy^2 + 2y \sqrt{x^2y^2+z} )dx + ( -2x^2y+ 2x \sqrt{x^2y^2+z} )dy = dz"

Integrating term by term, we have

"z = -x^2y^2 +xy\sqrt{x^2y^2 +z} + z ln|y(\sqrt{x^2y^2 +z}+xy)|+ a -x^2y^2 +xy\sqrt{x^2y^2 +z} + z ln|x(\sqrt{x^2y^2 +z}+xy)| + b"

i.e

"z = -2x^2y^2 +2xy\sqrt{x^2y^2+ z} + z [ln|x(\sqrt{x^2y^2+ z} + xy)| + ln|y(\sqrt{x^2y^2+ z} + xy)|] + a + b" where a and b are constants